### v2

@@ 1 Introduction
he rapid advancement of large vision language models (LVLMs) Xie et al. (2024); Liu et al. (2023, 2024); Dai et al. (2023); Zhuang et al. (2025); Chen et al. (2024); Yao et al. (2024); Weihan Wang and Qingsong Lv and Wenmeng Yu and Wenyi Hong and Ji Qi and Yan Wang and Junhui Ji and Zhuoyi Yang and Lei Zhao and Xixuan Song and Jiazheng Xu and Bin Xu and Juanzi Li and Yuxiao Dong and Ming Ding and Jie Tang (2024) has significantly enhanced the capabilities of multimodal agents, enabling them to process and integrate heterogeneous information sources such as visual scenes, natural language instructions, and structured external data. These multimodal agents have demonstrated impressive performance across a wide range of applications, including virtual assistants Pei et al. (2024); Chu et al. (2023); Hong, Wenyi and Wang, Weihan and Lv, Qingsong and Xu, Jiazheng and Yu, Wenmeng and Ji, Junhui and Wang, Yan and Wang, Zihan and Dong, Yuxiao and Ding, Ming and Tang, Jie (2024), autonomous driving systems Wang et al. (2023); Ma et al. (2024); Xu et al. (2024), and embodied agents Abdin et al. (2024); Wallace et al. (2024); Mu et al. (2023); Li et al. (2024). By leveraging the complementary strengths of different modalities, these systems can make more informed and context-sensitive decisions.
@@ 2.1 Multimodal Agent
As multimodal agents gain widespread adoption across critical domains, significant security vulnerabilities have emerged Liu et al. (2024); Wang et al. (2025); Zhang et al. (2025); Wu et al. (2024); Aichberger et al. (2025); Liu et al. (2020); Liang et al. (2020); Wei et al. (2018); Liang et al. (2022a, b); Yan et al. (2024); Ying and Wu (2023a, b). In this paper, we investigate the security implications Zhang et al. (2024); Kong et al. (2024); Ying et al. (2024, 2025); Liang et al. (2023, 2024a, 2024b); Ying et al. (2024a, b, 2025); Jing et al. (2025) of multimodal agents under cross-modal prompt injection attacks, with particular emphasis on their susceptibility to malicious inputs originating from diverse sources.
@@ 4 Methodology
Figure 2: Overall Framework. CrossInject implements a novel cross-modal prompt injection attack on multimodal agent systems. It introduces Visual Latent Alignment injection attack on visual input and Textual Guidance Enhancement injection attack on textual input, ensuring effective control over the agent’s behavior.
@@ 4.2 Textual Guidance Enhancement
Due to the attacker’s limited black-box access to the agent system, we consider inferring the defense-aware system prompt $S$ aligned with the agent’s safety mechanism. Here we employ a powerful language model (i.e., GPT-4 OpenAI (2023)) to infer safety-aligned system prompt through role-driven analysis. Typically, to amplify adversarial impact of textual injection, we assume strengthened defensive protection of the victim agent when inferring system prompt, pushing the textual optimization to discover the command $C^{\prime}$ that overcomes such strong protection. This adversarially leads to malicious command $C^{\prime}$ with enhanced control over the victim agent’s output. Specifically, we apply meta prompting Yifan Zhang and Yang Yuan and Andrew Chi-Chih Yao (2025) method to construct the inference prompts for different agents. Meta prompt is an example-agnostic structured prompt, which can be used to efficiently guide LLM to infer system prompts for agents with diverse roles. Let $R$ be denoted as role description of a victim agent, which is known to the attacker in our threat model, and let $D(\cdot)$ represent the defense-aware meta prompt template, the inferred system prompt $S^{\prime}$ is formulated as below:
@@ 5.1 Experimental Setup
Tested VLMs. We adopt two state-of-the-art VLMs, Qwen2-VL Wang et al. (2024) and Phi-3.5-vision Abdin et al. (2024), as the core planners for the agents. Qwen2-VL is a novel vision-language model that demonstrates superior performance across various visual understanding benchmarks, which is well suited for integration into agent systems, such as mobile phones and physical robots Team (2024). Phi-3.5-vision, developed by Microsoft, achieves cutting-edge performance in image understanding while maintaining robust safety capabilities Ying et al. (2024). For both models, we set the max new tokens parameter to 1024.
@@
Implementation Details. We set the visual perturbation budget $\epsilon$ to 16 under the $\ell_{\infty}$ constraint, a value commonly used in visual adversarial examples Carlini et al. (2019), and the number of iterations to 200. We employ four types of multimodal vision encoders as surrogate models for visual attacks, including ViT-H-14, ViT-L-14, ViT-B-16, and ViT-SO400M-14-SigLIP-384. For textual attacks based on GCG algorithm, we set the top-k value to 256, batch size to 512, and the number of iterations to 100. Specifically, we leverage Llama-3.1-8B-Instruct as the surrogate large language model for optimizing adversarial textual commands. The LLM judge employs Qwen-Max Team (2024), a powerful large language model. To ensure statistical reliability, each experiment was repeated three times, and the average results are reported. All code is implemented in PyTorch, and all experiments are conducted on an NVIDIA A800-SXM4-80GB GPU cluster.
@@ 5.3 Ablation Studies
Surrogate LLM for Textual Optimization. To further validate the role of surrogate LLMs in textual optimization, we additionally selected three open-source LLMs: Llama-2-7B, Vicuna-7B The Vicuna Team (2023) finetuned on Llama-2-7B, and Mistral-7B Mistral AI team (2023) based on Mixture of Experts (MoE) Noam Shazeer and *Azalia Mirhoseini and *Krzysztof Maziarz and Andy Davis and Quoc Le and Geoffrey Hinton and Jeff Dean (2017) architecture.

### v3

@@ 1 Introduction
he rapid advancement of large vision language models (LVLMs) Xie et al. (2024); Liu et al. (2023, 2024); Dai et al. (2023); Zhuang et al. (2025); Chen et al. (2024); Yao et al. (2024); Wang et al. (2024) has significantly enhanced the capabilities of multimodal agents, enabling them to process and integrate heterogeneous information sources such as visual scenes, natural language instructions, and structured external data. These multimodal agents have demonstrated impressive performance across a wide range of applications, including virtual assistants Pei et al. (2024); Chu et al. (2023); Hong et al. (2024), autonomous driving systems Wang et al. (2023); Ma et al. (2024); Xu et al. (2024), and embodied agents Abdin et al. (2024); Wallace et al. (2024); Mu et al. (2023); Li et al. (2024); Qin et al. (2023). By leveraging the complementary strengths of different modalities, these systems can make more informed and context-sensitive decisions.
@@ 2.1 Multimodal Agent
As multimodal agents gain widespread adoption across critical domains, significant security vulnerabilities have emerged Liu et al. (2024); Wang et al. (2025); Zhang et al. (2025); Wu et al. (2024); Aichberger et al. (2025); Liu et al. (2020); Liang et al. (2020); Wei et al. (2018); Liang et al. (2022a, b); Yan et al. (2024); Ying and Wu (2023a, b). In this paper, we investigate the security implications Zhang et al. (2024); Kong et al. (2024); ying2024safebench; Ying et al. (2025); Liang et al. (2023, 2024a, 2024b); Ying et al. (2024a, b, 2025); Jing et al. (2025) of multimodal agents under cross-modal prompt injection attacks, with particular emphasis on their susceptibility to malicious inputs originating from diverse sources.
@@ 4 Methodology
Figure 2: Overall Framework. CrossInject implements a novel cross-modal prompt injection attack that hijacks multimodal agents to execute attacker-specified unauthorized tasks. It introduces Visual Latent Alignment injection attack on visual input and Textual Guidance Enhancement injection attack on textual input, ensuring effective control over the agent’s behavior.
@@ 4.2 Textual Guidance Enhancement
Due to the attacker’s limited black-box access to the agent system, we consider inferring the defense-aware system prompt $S$ aligned with the agent’s safety mechanism. Here we employ a powerful language model (i.e., GPT-4 OpenAI (2023)) to infer safety-aligned system prompt through role-driven analysis. Typically, to amplify adversarial impact of textual injection, we assume strengthened defensive protection of the victim agent when inferring system prompt, pushing the textual optimization to discover the command $C^{\prime}$ that overcomes such strong protection. This adversarially leads to malicious command $C^{\prime}$ with enhanced control over the victim agent’s output. Specifically, we apply meta prompting Zhang et al. (2025) method to construct the inference prompts for different agents. Meta prompt is an example-agnostic structured prompt, which can be used to efficiently guide LLM to infer system prompts for agents with diverse roles. Let $R$ be denoted as role description of a victim agent, which is known to the attacker in our threat model, and let $D(\cdot)$ represent the defense-aware meta prompt template, the inferred system prompt $S^{\prime}$ is formulated as below:
@@ 5.1 Experimental Setup
Tested VLMs. We adopt two state-of-the-art VLMs, Qwen2-VL Wang et al. (2024) and Phi-3.5-vision Abdin et al. (2024), as the core planners for the agents. Qwen2-VL is a novel vision-language model that demonstrates superior performance across various visual understanding benchmarks, which is well suited for integration into agent systems, such as mobile phones and physical robots Qwen Team (2024). Phi-3.5-vision, developed by Microsoft, achieves cutting-edge performance in image understanding while maintaining robust safety capabilities Ying et al. (2024). For both models, we set the max new tokens parameter to 1024.
@@
Implementation Details. We set the visual perturbation budget $\epsilon$ to 16 under the $\ell_{\infty}$ constraint, a value commonly used in visual adversarial examples Carlini et al. (2019), and the number of iterations to 200. We employ four types of multimodal vision encoders as surrogate models for visual attacks, including ViT-H-14, ViT-L-14, ViT-B-16, and ViT-SO400M-14-SigLIP-384. For textual attacks based on GCG algorithm, we set the top-k value to 256, batch size to 512, and the number of iterations to 100. Specifically, we leverage Llama-3.1-8B-Instruct as the surrogate large language model for optimizing adversarial textual commands. The LLM judge employs Qwen-Max Qwen Team (2024), a powerful large language model. To ensure statistical reliability, each experiment was repeated three times, and the average results are reported. All code is implemented in PyTorch, and all experiments are conducted on an NVIDIA A800-SXM4-80GB GPU cluster.
@@ 5.3 Ablation Studies
Surrogate LLM for Textual Optimization. To further validate the role of surrogate LLMs in textual optimization, we additionally selected three open-source LLMs: Llama-2-7B, Vicuna-7B The Vicuna Team (2023) finetuned on Llama-2-7B, and Mistral-7B Mistral AI team (2023) based on Mixture of Experts (MoE) Shazeer et al. (2017) architecture.